In its recent year-end, Rod Roy reported an after tax loss of $10,000 which included income from continuing operations of $50,000 after tax. There were discontinued operations during the year.

IT also reported the following at the beginning of the year:

Preferred shares, 3%, cumulative                       $200,000

Common shares, 5,000 issued and outstanding   $100,000

No shares were issued or repurchased during the year, however there were outstanding options issued several years prior allowing for the purchase of 6,000 common shares at a price of $15. The average market price in the year was $18.

Required:

Provide the required IFRS disclosures for earnings per share data required in Rod Roy’s annual report.
Answer #1

Profit after Tax= $ 50000

Dividend to Pref. Share = $200000*3% = $6000

Profit after Dividend to Pref Share holder = $(50000-6000) = $44000

Earning Per Share = $44000/5000 = $8.8

Diluted Earing Per Share = $44000/(5000+6000) = $44000/11000 = $4